# No Vulnerability Data, No Problem: Towards Predicting Mean Time To Remediate In Open Source Software Dependencies

### v2

@@ 6.2. RQ1: How do other package characteristics influence MTTUdep and MTTRdep?
Comparison to Prior Work. Similar to prior research (kula_developers_2018, ), we also found that the majority of packages in an ecosystem do not keep dependencies updated. We also observed a long-tail distribution for both MTTUdep and MTTRdep in all three ecosystems. A significant number of packages lag behind in keeping dependencies updated, which is similar to the observations of Cox et al. (cox_measuring_2015, ). In short, regularly updating dependencies is not a widespread practice. Our observation is similar to vulnerable dependency updates as well. Many packages took months or longer to remediate known vulnerable dependencies, similar to observations from Kula et al. (kula_developers_2018, ). This suggests that challenges in keeping dependencies updated are not specific to any ecosystem but rather general in nature.
@@ 7. Discussions And Implications
When MTTUdep and MTTRdep should be used. Our study provides new insights by establishing the relationship between MTTUdep and MTTRdep. We found that although a package’s MTTUdep is moderately correlated with its MTTRdep, MTTUdep and MTTRdep are statistically equivalence. In practical terms, packages that are slow to update dependencies also tend to be slow in updating vulnerable dependencies. This finding implies that improving general dependency update practice will likely improve vulnerable dependency update practice, which seems logical, but we provide statistical evidence supporting our hypothesis. Having said that, developers should use MTTRdep when available and only MTTUdep as a proxy for MTTRdep when MTTRdep is not available.
Package Age’s Influence In MTTUdep and MTTRdep. Package age is weakly correlated with MTTUdep and has a negligible correlation with MTTRdep in our analysis. In our regression analysis, we have used package age as an independent variable with MTTUdep and MTTRdep as dependent variables. However, since we design our metrics based on a temporal perspective, our metrics could have an inherent age sensitivity. So, conclusions like ‘older packages tend to have slightly higher MTTUdep’ should be treated cautiously. Since older packages might have more data points (negligible but positive correlation with version count, major version count, and dependencies count), package age could influence our analysis. Future research should explore how to model the package age in MTTUdep and MTTRdep (e.g., one way could be to normalize the metrics by package age or by only taking the last $n$ years into account).
Influence of Unmodeled Package Characteristics. Prior work shows that developers hesitate to update dependencies due to a fear of breaking changes, a lack of awareness or knowledge about available updates, and sometimes a lack of motivation to invest time in updates (derrKeepMeUpdated2017, ). These human factors are the likely reasons behind the long tail of MTTUdep and MTTRdep data. Our findings also support this. Packages with many contributors or higher dependents count do not have a better MTTUdep or MTTRdep, in contrast to our hypothesis. A package used by thousands of other packages is not guaranteed to have dependencies updated. Both multilinear regression models in our study present an R-squared value of less than 10%. A lower R-squared value indicates that other unmodeled factors substantially affect MTTUdep and MTTRdep. The development team dynamic, the use of pinning and floating, the cost and efforts needed in testing, and the size of the codebase could be such possible unmodeled factors.
@@
Practical implications for developers. We confirmed that improving general dependency update frequency likely improves security as well. This should be a strong incentive for the developers to treat dependency updates as an important part of routine maintenance, not an afterthought. Even scheduling periodic dependency update sprints or using automated update notifications to stay on top of the new releases can be an effective strategy. Allowing the latest version of the dependency using the dependency specification (e.g., * or $latest$) is the best way to keep MTTUdep and MTTRdep low (even zero). However, that might not be possible due to the issues with breaking changes (ochoa_breaking_2022, ; raemaekers_semantic_2014, ; venturini_i_2023, ). Allowing auto updates of minor and patch releases by the dependency could help reduce the MTTUdep and MTTRdep. This ensures developers get incremental improvements and fixes without manual effort. Similarly, pinning should be avoided since pinning does not allow any auto-updates, which will start increasing the MTTUdep as soon as the dependency releases a new version (even a patch release). Whenever a vulnerability is found in the dependency and a fixed version is available, developers should prioritize remediating that vulnerability. Finally, a mixed strategy with automatic minor and patch updates, alongside manual major version updates, could be the most effective strategy to keep the MTTUdep and MTTRdep minimal. Our recommendation relies on developers using SEMVER correctly (pinckney_large_2023, ; li_large-scale_2023, ; dietrich_dependency_2019, ) so that the benefits of floating versions can be leveraged.
@@
Gaming Metrics. While it might seem that using a loose floating constraint and bumping dependency constraint periodically could “game” our metrics, the underlying outcome is that a package spends minimal time with outdated dependency versions and ensures that security fixes are adopted automatically. By doing so, this strategy reduces the risk of exploitation from using an outdated dependency version with known security vulnerability, such as log4Shell (everson_log4shell_2022, ; rahman_less_2024, ). In short, trying to game our metrics by adopting floating version constraints is essentially reducing this attack vector for a package. However, this approach also comes with a tradeoff. This strategy might make the packages susceptible to malicious package updates (xz-incident (williams_research_2025, )). Since floating version constraints allow automatic updates, it can also facilitate propagating malicious package updates.
Influence of Dependencies Count in MTTUdep and MTTRdep. In our design, we compute MTTUdep and MTTRdep by dividing the summation of TTUdep and TTRdep by the number of dependencies and number of historically vulnerable dependencies, respectively, in Eq. 3 and  5. A package with a high number of dependencies has more room for having a better value in MTTUdep while keeping the majority of dependencies updated. However, MTTRdep is not influenced by dependencies count since we only consider the amount of time spent with vulnerable dependency versions and the number of historically vulnerable dependencies.
@@ 8. Threats to Validity
In this section, we discuss potential limitations and how they might impact the interpretation of our results.
External Validity. The main external validity threat is the generalizability of our results to characterize other ecosystems. We carefully chose the three ecosystems to have a diverse set of ecosystems, where npm is the largest, PyPI is the oldest, and Cargo is the newest among the major software ecosystems. While each ecosystem possesses unique features that might not directly correlate with those we studied, we believe the insights gained should also be broadly applicable to other ecosystems.
Internal Validity. We use the security advisory dataset from OSV.dev, which may not be comprehensive. If an advisory is published but not included in the OSV dataset, that may impact our results. Additionally, we do not consider whether the vulnerable dependency version is exploitable or reachable (cisa-vex, ) from the package. We treat all vulnerabilities equally regardless of the CVSS score or the severity of the vulnerability. After downloading the data from deps.dev, we manually checked 20 packages’ versions and relations with the public package registries and found that the data is accurate. Moreover, each package manager has its own way of handling dependency resolutions, and for the dependency resolutions, we rely on the Open Source Insights (depsdev, ) resolved version data. Our analysis omits package versions not adhering to SEMVER rules, a conservative choice to enable a more rigorous analysis. In addition, Open Source Insights dependency resolution fails in some cases (e.g. missing timestamp), and we mark those as warnings in our dataset. We do not calculate update metrics for those packages, and we argue that this might have a very small impact on our results. In addition, we only consider runtime dependencies in our MTTUdep and MTTRdep analysis and omit dev and optional dependencies. Additionally, some dependencies might be more important than others depending on the context; however, we treat each dependency equally since modeling dependencies’ importance is out of the scope of our study.

### v3

@@ 6.1. RQRQ2:: : How do packages in npm, PyPI, and Cargo differ in MTTU and MTTR?
Comparison to Prior Work. In contrast to prior research (kula_developers_2018, ), we found that the majority of packages in an ecosystem have a lower MTTRdep. The reason behind our metrics encompassing lower MTTRdep values is that our metrics weight recency, and older dependency update practice has less impact on our metrics. We also observed a long-tail distribution for both MTTUdep and MTTRdep in all three ecosystems. A significant number of packages lag behind in keeping dependencies updated, which is similar to the observations of Cox et al. (cox_measuring_2015, ). In short, regularly updating dependencies might not be a widespread practice. Our observation is similar to vulnerable dependency updates as well. Even with weighting, some packages took months or longer to remediate known vulnerable dependencies, similar to observations from Kula et al. (kula_developers_2018, ). This suggests that challenges in keeping dependencies updated are not specific to any ecosystem but rather general in nature.
Takeaway 2: Most packages in npm, PyPI, and Cargo have relatively fast dependency update practices. The small interquartile ranges indicate consistent dependency update practice with each ecosystem.
Table 3. TOSTs Results
Ecosystem MTTUdep $\times 0.8$ << MTTRdep << MTTUdep $\div 0.8$
All $U=$2.5\text{.}{10}^{08}$\$, $p=0.02$ $U=$1.9\text{.}{10}^{08}$\$, $p<0.01$
Cargo $U=4668$, $p=0.51$ $U=2806$, $p=$7.8\text{.}{10}^{-5}$\$
npm $U=$1.5\text{.}{10}^{08}$\$, $p=0.04$ $U=$1.2\text{.}{10}^{08}$\$, $p=$6.1\text{.}{10}^{-278}$\$
PyPI $U=$1.3\text{.}{10}^{07}$\$, $p=0.0005$ $U=$1.0\text{.}{10}^{07}$\$, $p=$5.6\text{.}{10}^{-103}$\$
Table 4. Proxy Analysis Result
Criteria Result Pass?
TOST Statistically significant Yes
Regression - Coefficient Statistically significant Yes
Regression - $R^{2}$ Moderate Moderately
Regression - Wald test Statistically significant Yess
Sensitivity - GFI <0.90<0.90 No
Sensitivity - AGFI <0.90<0.90 No
Sensitivity - NFI <0.90<0.90 No
Correlation - Pearson Moderate and positive Moderately
Correlation - Spearman Moderate and positive Moderately
6.2. RQRQ3:: : Can MTTU serve as a proxy for MTTR?
As we specified in Section 5.4, the proxy analysis covered four tests/ analyses:
(1) TOST (Two One-Sided Test). The results of the TOSTs are reported in Table 4. The results of the TOSTs show statistical equivalence between MTTUdep and MTTRdep ($p<0.05$), except for Cargo ($p>0.05$).
(2) Regression with Wald Test. The coefficient of MTTUdep from the OLS regression returned positive and also statistically significant ($\beta=0.69,SE=0.056,t=12.21,p<0.001$), which indicates that, on average, when MTTUdep increases by one unit, MTTRdep also increases by 0.69. The $R^{2}$, however, shows that MTTUdep only explained 30.5% of the MTTRdep’s variance ($R^{2}=0.305$). The interpretation of this $R^{2}$ varies in different domains. If we refer to Hair et al. (hair2022partial, ), we can interpret this $R^{2}$ as a moderate explanatory power. The Wald test also returns statistically significant ($F=149.1,p<0.001$), which confirms that MTTUdep contributes to explaining MTTRdep.
(3) Sensitivity Analysis. The sensitivity analysis returns $GFI=0.83,AGFI=0.67$, and $NFI=0.83$. As these indices are lower than the acceptable threshold (0.90), this result indicates that MTTUdep does not provide a good fit to MTTRdep.
(4) Correlation Analysis. Pearson correlation coefficients returns 0.552 with $p=$1.12\text{.}{10}^{-28}$\$. Spearman, on the other hand, returns 0.689 with $p=$1.47\text{.}{10}^{-49}$\$. The correlations show that MTTUdep and MTTRdep are moderately and positively correlated. This correlation shows that MTTUdep scales linearly with MTTRdep.
The results from the four tests/analyses are summarized in Table 4. In summary, (1) MTTUdep is statistically equivalent to MTTRdep; (2) MTTUdep increases along with the increases of MTTRdep, with a moderate explanatory power; (3) MTTUdep does not provide an acceptable fit for MTTRdep; and (4) MTTUdep is moderately and positively correlated with MTTRdep. Considering all metrics as distinct criteria, MTTUdep strongly satisfied three criteria, moderately satisfied three others, and failed to meet the other three left, suggesting partial adequacy as a proxy for MTTRdep. This suggested partial adequacy indicates that developers may use MTTUdep as a proxy for MTTRdep with caution and encourages further research to find better proxies for MTTRdep.
Takeaway 3: MTTUdep can only partially serve as a proxy for MTTRdep. This suggests that MTTUdep may be used as a proxy for MTTRdep with caution, and future research is needed to find a better proxy.
6.3. RQRQ4:: : How do package characteristics influence MTTU and MTTR?
Correlation Analysis. We present the correlation matrix with the nine packages’ characteristics and MTTUdep and MTTRdep data in Figure 4. The correlation heatmap provides insight into the relationships between various factors and their potential impact on MTTUdep and MTTRdep. The results indicate weak correlations between MTTUdep and the other factors.
Contributors Count ($H_{1}$): We hypothesized that packages with fewer contributors would be less likely to maintain updated dependencies. From the heatmap, we can see that the contributors count shows minimal correlations with MTTUdep (-0.07) and MTTRdep (-0.07), suggesting a limited influence on dependency management.
Dependents Count ($H_{2}$): We hypothesized that packages with fewer dependents are less likely to have updated dependencies. The results indicate a weak negative correlation between dependents count with MTTUdep (-0.17) and MTTRdep (-0.16). This suggests that the number of dependents plays a limited role in determining the timeliness of dependency updates.
Dependency Count ($H_{3}$): We hypothesized that packages with fewer dependencies would have more updated dependencies. The results show a weak negative correlation between dependency count with MTTUdep (-0.10) and MTTRdep (-0.11). This implies that the number of dependencies in a package has little association with dependency management efficiency.
Version Count ($H_{4}$): We hypothesized that packages with fewer version releases are less likely to have updated dependencies. However, the correlation matrix revealed a moderate negative relationship between the number of version releases and MTTUdep (-0.55) and MTTRdep (-0.41). This supports our hypothesis that packages with more version releases are likely to have updated dependencies.
Major Version Count ($H_{5}$): We hypothesized that packages with fewer major version releases are more likely to have updated dependencies. However, the correlation matrix revealed a weak negative relationship between the number of major version releases and MTTUdep (-0.16) and MTTRdep (-0.14). This suggests that the number of available major versions to maintain has a limited impact on how quickly dependencies are updated.
Package Age ($H_{6}$): We hypothesized that packages with lower ages are more likely to have updated dependencies. However, the correlation matrix reveals MTTUdep (0.10) and MTTRdep (0.07) have a weak correlation with package age. This suggests that package age has little impact on having updated dependencies. This also strengthens our intuition of creating weighted versions, MTTUdep and MTTRdep, to eliminate the potential age-sensitivity of MTTUdep and MTTRdep.
SourceRank ($H_{7})$. We hypothesized that low quality packages (e.g., packages with lower SourceRank scores) are less likely to have updated dependencies. The correlation matrix shows limited support for this hypothesis since low-quality packages do not necessarily have weak correlation with MTTUdep (0.09-0.09) and MTTRdep (0.10-0.10).
Forks Count ($H_{8}$): We hypothesized that packages with fewer forks are less likely to have updated dependencies. The results reveal a weak to moderate positive correlation between forks (0.86) with maintainers count, indicating that more forked packages may have more active development. However, the relationships between forks count with MTTUdep and MTTRdep remain insignificant, suggesting that having more forks does not directly translate into faster dependency updates.
Stars Count ($H_{9}$): We hypothesized that packages with fewer stars are less likely to have updated dependencies. There is a weak to moderate positive correlation between stars and maintainers count (0.85), indicating that more popular packages may have more active development. However, the relationships between star count with MTTUdep and MTTRdep are insignificant, implying that more stars do not directly indicate faster dependency updates.
Overall, our results indicate weak or negligible correlations between the tested factors and MTTUdep and MTTRdep. These findings suggest that the analyzed factors may not strongly influence how updated dependencies are, and further exploration of other variables or non-linear relationships is recommended.
Takeaway 4: Package characteristics (except version count) have negligible correlations with dependency updatedness. Packages with higher version count are associated with lower MTTUdep and MTTRdep.
Regression Analysis (MTTUdep As Dependent Variable). Interpreting the multilinear regression model’s characteristics, we found 0.039 as the $R^{2}$ value when MTTUdep was the dependent variable. $R^{2}$ value 0.039 indicates that 3.9% variation in the dependent variable can be explained by the model. Although the model is statistically significant ($p_{value}$ ¡ 0.0055 from the F-test), a lower $R^{2}$ value indicates that other factors, beyond the nine dependent variables, also substantially impact MTTUdep. The F-statistic of this model is large (547.7), and Prob(F-statistic) is 0.00. This indicates that at least one of the independent variables (or predictors) has a non-zero relationship with MTTUdep.
We then look into the coefficients and $p_{value}$s associated with each of the independent variables. The coefficient indicates the expected change in the dependent variable for a one-unit change in one independent variable while holding other independent variables constant. The $p_{value}$ associated with the t-statistic indicates whether the independent variable is statistically significant in explaining the variation in the dependent variable. A lower $p_{value}$ (¡ 0.0055) would indicate statistical significance.
We have found that some independent variables have positive coefficients, which indicates that an increase in one of these variables would result in an increase in our dependent variable, MTTUdep. Independent variables with positive coefficients are contributors count (0.0025), dependents count ($5.3\text{.}{10}^{05}$), forks (0.0002), and package age (0.0114). The package age has the largest positive coefficient, which indicates older packages tend to have longer MTTUdep, justifying our hypothesis ($H_{6}$) on package age. However, the rest of the coefficients are small, which indicates these independent variables do not have a significant impact on MTTUdep.
We also found that some independent variables have negative coefficients. Negative coefficients indicate that an increase in that independent variable would result in a decrease in the dependent variable, MTTUdep. Independent variables with negative coefficients are SourceRank (-2.2260), stars (-0.0002), dependencies count (-0.0493), and major versions count (-0.4766). SourceRank, which indicates package quality and popularity, shows moderate negative coefficients. Higher SourceRank indicates an overall better project. The model predicts that if SourceRank for a package increases by 1 (moving to a “better” rank), MTTUdep drops by 2 days. Moreover, packages that have released more major versions tend to have a lower MTTUdep. Finally, independent variables, except dependent count and forks, were found to be statistically significant ($p_{value}$ ¡ 0.0055).
Takeaway 5: Older packages are more likely to have higher MTTUdep. Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTUdep.
Regression Analysis (MTTRdep As Dependent Variable). For MTTRdep as the dependent variable, the multilinear regression model results in 0.026 as the $R^{2}$ value. $R^{2}$ = 0.026 indicates that the model explains 2.6% of the variation in MTTRdep. A lower $R^{2}$ value indicates that other unmodeled factors play an important role in determining MTTRdep, which is similar to our observations in MTTUdep. The F-statistic of this model is 65.4, and Prob(F-statistic) is $2.5\text{.}{10}^{-119}$. This indicates the model is statistically significant and at least one of the dependent variables (or predictors) has a non-zero relationship with MTTRdep.
We then look into the coefficients and $p_{value}$s associated with each of the independent variables. We found that independent variables, except for the dependent count, forks, and stars, are statistically significant ($p_{value}$ ¡ 0.0055) in modeling MTTRdep.
As a predictor, forks (coefficient 0.0008), dependents count (coefficient $2.3\text{.}{10}^{-05}$ and package age (coefficient 0.0091) indicate a positive but small effect. Older packages are associated with slightly higher MTTRdep. Additionally, the major version count (coefficient -0.3453), dependencies count (coefficient -0.1213), contributors count (coefficient -0.0107), and SourceRank (coefficient -1.3284) show a moderate negative effect on MTTRdep. This observation is similar to our previous observation with MTTUdep. According to the model’s prediction, an increase in SourceRank by 1 (i.e., a higher SourceRank indicates a better package) would result in a 1-day reduction in MTTRdep.
Takeaway 6: Similar to MTTUdep, older packages are more likely to have higher MTTRdep. Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTRdep. Packages with higher major versions are more likely to have lower MTTRdep.
@@ 7.1. Practical Implications For Developers
When MTTUdep And MTTRdep Should Be Used. Our study provides new insights by exploring the relationship between MTTUdep and MTTRdep. While MTTUdep does not fully meet all the criteria to serve as a proxy for MTTRdep, it satisfies six out of nine criteria, which can be considered as a partial proxy. This finding provides enough evidence for MTTU to be a practical indication of MTTR, especially when externally reported vulnerability data is not available. In practical terms, packages that are slow to update dependencies also tend to be slow in updating vulnerable dependencies. This finding implies that improving general dependency update practice will likely improve vulnerable dependency update practice as well. Having said that, developers should use MTTRdep when available and only MTTUdep as a proxy for MTTRdep with caution when MTTRdep is not available.
@@
Use Floating-Minor With Regular Major Updates. Our results indicate that improving general dependency update frequency likely improves security as well. This should be a strong incentive for the developers to treat dependency updates as an important part of routine maintenance, not an afterthought. Even scheduling periodic dependency update sprints or using automated update notifications to stay on top of the new releases can be an effective strategy. Allowing the latest version of the dependency using the dependency specification (e.g., * or latestlatest) is the best way to keep MTTUdep and MTTRdep low (even zero). However, that might not be possible due to the issues with breaking changes (ochoa_breaking_2022, ; raemaekers_semantic_2014, ; venturini_i_2023, ). Allowing auto updates of minor and patch releases by the dependency could help reduce the MTTUdep and MTTRdep. This ensures developers get incremental improvements and fixes without manual effort. Similarly, pinning should be avoided since pinning does not allow any auto-updates, which will start increasing the MTTUdep as soon as the dependency releases a new version (even a patch release). Whenever a vulnerability is found in the dependency and a fixed version is available, developers should prioritize remediating that vulnerability. Finally, a mixed strategy with automatic minor and patch updates, alongside manual major version updates, could be the most effective strategy to keep the MTTUdep and MTTRdep minimal. Our recommendation relies on developers using SEMVER correctly (pinckney_large_2023, ; li_large-scale_2023, ; dietrich_dependency_2019, ) so that the benefits of floating versions can be leveraged.
Dependency Selection Criteria. Our findings suggest that packages with higher SourceRank scores showed somewhat more efficient update practice. In practice, this means a popular, actively maintained package is more likely to receive timely updates (and even external contributions) than an obscure one. When choosing a dependency, a well-maintained package should be prioritized if that serves the required functionality. In addition, our findings suggest that packages with a higher number of versions and a higher number of major versions are likely to have better practices for updating vulnerable dependencies. A higher number of versions and major versions indicates the development team of the package has released additional features, is active in maintenance, and is more responsive to making changes. In case of a vulnerability in dependencies, this development team might be more prompt in releasing a new version of the package with mitigating the vulnerability, or they use a version constraint (e.g., floating) which automatically adopts security fixes from their dependencies.
7.2. Practical Implication For Researchers
Prior work shows that developers hesitate to update dependencies due to a fear of breaking changes, a lack of awareness or knowledge about available updates, and sometimes a lack of motivation to invest time in updates (derrKeepMeUpdated2017, ). These human factors are the likely reasons behind the long tail of MTTUdep and MTTRdep data. Our findings also support prior research, revealing that packages with many contributors or higher dependents count do not have a better MTTUdep or MTTRdep, in contrast to our hypothesis. Even packages with thousands of dependents are not guaranteed timely updates for their own dependencies. Researchers should focus more on exploring the human factors to uncover ways to balance the cost and benefit of dependency updates. In addition, both multilinear regression models in our study present an $R^{2}$ value of less than 10%. A lower $R^{2}$ value indicates that other unmodeled factors substantially affect MTTUdep and MTTRdep. The development team dynamics, the use of pinning and floating, the cost and efforts needed in testing, and the size of the codebase could be such possible unmodeled factors. Researchers should explore further if such other unmodeled factors influence MTTUdep and MTTRdep.
7.3. Practical Implication For Tool Builders
In this study, we provide an extensive evaluation with our proposed novel dependency update metrics. Tool builders can incorporate our dependency update metrics into dependency management tools for developers to make them more accessible. Wermke et al. (wermkeAlwaysContributeBack2023, ) found that developers’ dependency selection metrics and criteria focus on quickly accessible numbers and facts, such as downloads, GitHub stars, and time since last release, substantiating the necessity of having easily comprehensible metrics. To this end, our metrics give quickly accessible a single number for a package that is quickly accessible and actionable for developers. Similarly, security risk assessment tools (e.g., OpenSSF Scorecard) could also incorporate our dependency update metrics in their assessment to help developers better assess the security risk of a package.
7.4. Gaming Metrics
@@
While it might seem that using a loose floating constraint and bumping dependency constraint periodically could “game” our metrics, the underlying outcome is that a package spends minimal time with outdated dependency versions and ensures that security fixes are adopted automatically. By doing so, this strategy reduces the risk of exploitation from using an outdated dependency version with a known security vulnerability, such as log4Shell (everson_log4shell_2022, ; rahman_less_2024, ). In short, trying to game our metrics by adopting floating version constraints is essentially reducing this attack vector for a package. However, this approach also comes with a tradeoff. This strategy might make the packages susceptible to malicious package updates (xz-incident (williams_research_2025, )). Since floating version constraints allow aucan also facilitate propagating malicious package updates.
@@ 8. Threats to Validity
External Validity. The main external validity threat is the generalizability of our results to characterize other ecosystems. While each ecosystem possesses unique features that might not directly correlate with those we studied, we believe the insights gained should also be broadly applicable to other ecosystems.
Internal Validity. We use the security advisory dataset from OSV.dev, which may not be comprehensive. If an advisory is published but not included in the OSV dataset, that may impact our results. Additionally, we do not consider whether the vulnerable dependency version is exploitable or reachable (cisa-vex, ) from the package. We treat all vulnerabilities equally, regardless of the CVSS score or the severity of the vulnerability. Using the severity of vulnerabilities as a weighting factor in our metrics would be an interesting future work. After downloading the data from deps.dev, we manually checked 20 packages’ versions and relations with the public package registries and found that the data is accurate. Moreover, each package manager has its own way of handling dependency resolutions, and for the dependency resolutions, we rely on the Open Source Insights (depsdev, ) resolved version data. Our analysis omits package versions not adhering to SEMVER rules, a conservative choice to enable a more rigorous analysis. In addition, Open Source Insights dependency resolution fails in some cases (e.g., missing timestamp), and we mark those as warnings in our dataset. We do not calculate update metrics for those packages, and we argue that this might have a very small impact on our results. In addition, we only consider runtime dependencies in our MTTUdep and MTTRdep analysis and omit dev and optional dependencies. Additionally, some dependencies might be more important than others depending on the context; however, we treat each dependency equally since modeling dependencies’ importance is out of the scope of our study.